OpenStudy (anonymous):

You want to purchase a house. You need to borrow $175,000 and can find a 30 year loan that wants monthly payments and 6.24% interest compounded monthly. What is the size of your monthly payment?

OpenStudy (tkhunny):

Learn to to build it from scratch. i = 0.0624 -- Annual Interest Rate, to be compounded monthly. j = i/12 = 0.0624/12 = 0.0052 -- Monthly interest rate. v = 1/(1+j) = 0.994826900 -- Monthly Discount Factor \(175000 = Pmt(v + v^{2} + v^{3} + ... + v^{360})\) That's it. The rest is algebra. \(175000 = Pmt\cdot\dfrac{v - v^{361}}{1-v}\) \(175000\cdot \dfrac{1-v}{v-v^{361}} = Pmt\) You'll never wonder about the right formula again if you get the idea.

OpenStudy (anonymous):

so which formula do I use?

OpenStudy (anonymous):

@tkhunny so which formula do I use